A Bibliography of Publications of Ulrich Nitsche

Ulrich Nitsche University of Zurich

Department of Computer Science

Winterthurer Str. 190

CH-8057 Zurich Switzerland

Tel: ?n/a? FAX: ?n/a?

E-mail: nitsche@ifi.unizh.ch (Internet)

19 February 2013 Version 1.06

Abstract

This bibliography records publications of Ulrich Nitsche.

Title word cross-reference

∗ [Nit97]. ω [NNO97, NNO98].

-Regular [NNO97, NNO98].

/Telekom [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b].

31 [AH96]. 31-August [AH96].

4th [AN97].

8th [AH96].

’94 [NM94]. ’96 [AH96]. ’97 [AN97].

Abschlußbericht [DFVGE^{+}95, DFVGE^{+}96a]. Abstract [MNT97b, NW97]. Abstraction [CDGE^{+}96, CDFV^{+}96, Nit94c, Nit96b, NW97, Nit98, Nitxx]. Abstractions [Nit94d]. adequate [BAUN07]. aided [AH96]. Angeles [ACM94]. Annual [ACM94]. Application [Nit98]. Applications [Wag97]. Approach [BHM^{+}97]. Approximately [NO95a, NO96a, NO96b]. Approximation [UN02]. August [ACM94, ACM97, AH96]. Auswertung [Nit88]. automata [UN07].

Barbara [ACM97]. Based [Nit94c]. Behandlung [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Behavior [CDFV^{+}96, Nit96b, NW97]. Behaviour [CDGE^{+}96, Nit98, Nitxx]. biology [DRS96]. Brunswick [AH96]. Büchi [UN07].

California [ACM94, ACM97]. Care [MNT97c, MNT98]. Case [CDGE^{+}96, CDFV^{+}96]. CAV [AH96]. characterization [NUN07]. Co [Nitxx]. Co-Operating [Nitxx]. complete [NUN07]. Computation [Nit97]. Computer [AH96, AN97, DRS96, NM94]. Computing [ACM94, ACM97]. Concerns [MNT97c]. conference [AH96]. construction [UN07]. crossroads [DRS96]. CTL [Nit97].

Database [Wag97]. Datenschutz [MNT97a]. degree [UN07]. Detection [CDGE^{+}96, CDFV^{+}96]. determinism [UN07]. Deterministic [NNO97, NNO98, NUN07]. Developments [DRS96]. Distributed [ACM94, ACM97, Nit96b]. Do [UN02].

Efficiency [MNT97b, MNT98]. Eighth [Wag97]. eingeschränkte [Nit91a]. EIS [BAUN07]. Enterprise [AUN04]. Erreichbarkeitsanalyse [Nit91b]. Erreichbarkeitsgraphen [Nit88]. Erreichbarkeitsproblem [Nit91a]. Expert [Wag97]. Extended [NW97].

Finitary [Nit96a]. Focused [MNT98]. Formal [CDGE^{+}96, Nit98]. Formale [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Formulas [Nit94d]. foundations [AN97, NM94]. France [Wag97]. Full [Nit97].

Germany [DRS96]. GMD [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. GMD- [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. GMD-/Telekom-Projekts [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Graphs [Nit94d].

Health [MNT97c, MNT98]. Homomorphic [Nit94c]. Homomorphisms [Nit94a, Nit94b, NO95a, NO96a, NO96b]. ihre [Nit88]. II [DRS96]. Infinite [Nit97]. Information [AUN04, MNT97c]. Intelligent [CDFV^{+}96, Nit98]. Interaction [CDGE^{+}96, CDFV^{+}96, DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b, Nit98]. International [AUN04, Wag97, AN97, AH96, NM94]. Interpretation [Nit97].

July [AN97, AH96, DRS96, NM94].

Krankenhausinformationssysteme

[MNT97a].

Language [Nit94a, NO95a, NO96a, Nit96a, DRS96, NO96b]. Large [Nit94d, Nit96b]. LFCS [AN97, NM94]. Linear [Nit94a, Nit94b, Nit96a]. Liveness [NNO97, NNO98, NW97, UN02, NUN07]. Logic [Nit94a, Nit94b, Nit94d, Nit96a, Nit97]. Logical [AN97, NM94].

Magdeburg [DRS96]. mathematics [DRS96]. Medical [BHM^{+}97, MNT97b]. Method [Nit94c]. Mobile [MNT97b, MNT97c]. MobiMed [BHM^{+}97, MNT97b]. Model [Nit94c].

Need [UN02]. Networks

[CDFV^{+}96, Nit98]. NJ [AH96]. non [UN07]. non-determinism [UN07].

Operating [Nitxx].

Patient [MNT98]. Petersburg [NM94]. Poster [MNT97b]. power [UN07]. power-set [UN07]. Principles [ACM94, ACM97]. Privacy [BHM^{+}97, MNT97b, MNT98]. Problem [Nit98]. Problematik [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Proceedings [ACM94, ACM97, Wag97, AH96, NM94].

Processes [MNT98]. Produktnetze [Nit91a]. Produktnetzen [Nit88, Nit91b]. Projekts [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. PROLOG [Nit88]. Properties [NNO97, NNO98, NO95a, NO96a, UN02, NUN07, NO96b]. Propositional [Nit94a, Nit96a].

Qualitätskriterium [MNT97a].

Reachability [Nit94d]. reducing [UN07]. Regular [NNO97, NNO98, NUN07]. Relative [NW97]. Russia [AN97, NM94].

Safety [UN02]. Santa [ACM97]. Satisfied [NO95a, NO96a, NO96b]. science [AN97, DRS96, NM94]. Second [AUN04]. Security [MNT97c]. Semantics [Nit96a]. September [Wag97]. Sequences [Nit97]. Service [CDGE^{+}96, CDFV^{+}96, DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b, Nit98]. Service-Interaction-Problematik [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. SERVINT [DFVGE^{+}95, NO95b]. SERVINT2 [DFVGE^{+}96b]. SERVINTII [DFVGE^{+}96a]. Session [MNT97b]. set [UN07]. Sets [Nit97]. SH [NOR96, NORR97]. SH-Verification [NOR96, NORR97]. Simple [Nit94b, NO95a, NO96a, NO96b]. Sixteenth [ACM97]. Spezifikations [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Spezifikations- [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. St [NM94]. Study [CDGE^{+}96, CDFV^{+}96]. Symposium [ACM94, ACM97, AN97, NM94]. Systems [AUN04, BHM^{+}97, MNT97b, MNT97c, NO95a, NO96a, Nit96b, Nitxx, Wag97, NO96b].

Technique [Nit96b]. Telekom-Projekts [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Telephone [CDFV^{+}96].

Temporal [Nit94a, Nit94b, Nit94d, Nit96a]. theory [DRS96]. third [NM94]. Thirteenth [ACM94]. Tool [NORR97, NOR96]. Toulouse [Wag97]. Tractable [Nit96b]. Tree [Nit97]. two [UN07].

USA [ACM94, AH96].

Validation [AUN04]. Verification [AUN04, CDGE^{+}96, CDFV^{+}96, Nit94c, Nit96b, NORR97, Nit98, Nitxx, AH96, NOR96]. Verifikationsmethoden [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Verifying [Nit94d].

Workshop [AUN04, Wag97].

Yaroslavl [AN97].

zur [DFVGE^{+}95, DFVGE^{+}96a, DFVGE^{+}96b, NO95b]. Zwischenbericht [DFVGE^{+}96b, NO95b].

References

ACM:1994:PTA

[ACM94] ACM, editor. Proceedings of the Thirteenth Annual ACM Symposium on Principles of Distributed Computing: Los Angeles, California, USA, August 14–17, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-654-9. LCCN QA76.9.D5 A33 1994.

ACM:1997:PSA

[ACM97] ACM, editor. Proceedings of the Sixteenth ACM Symposium on Principles of Distributed Computing: Santa Barbara, California, August 21–24, 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-952-1. ACM order number 536970.

Alur:1996:CAV

[AH96] Rajeev Alur and T. A.

(Thomas A.) Henzinger, editors. Computer aided verification: 8th international conference, CAV ’96, New Brunswick, NJ, USA, July 31-August 3, 1996: proceedings, volume 1102 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61474-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1996.

Adian:1997:LFC

[AN97] Sergei Adian and Anil Nerode, editors. Logical foundations of computer science: 4th international symposium, LFCS ’97, Yaroslavl, Russia, July 6–12, 1997. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. LCCN QA75.5 .L43 1997.

Augusto:2004:SIW

[AUN04] J. C. Augusto and U. Ultes-Nitsche. Second international workshop on verification and validation of enterprise information systems. J.UCS: Journal of Universal Computer Science, 10(11):1495– 1497, November 28, 2004. CODEN ???? ISSN 0948-

                                        6968.  URL http://www.
                                        jucs.org/free;   http://
                                        www.jucs.org/jucs_10_11/
                                        second_international_workshop_
                                        on.

Barjis:2007:TMA

[BAUN07] Joseph Barjis, Juan Carlos Augusto, and Ulrich Ultes-Nitsche. Towards more adequate EIS. Science of Computer Programming, 65(1):1– 3, March 1, 2007. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bauknecht:1997:MAP

[BHM^{+}97] Kurt Bauknecht, Ralph Holbein, Othmar Morger, Ulrich Nitsche, and Stephanie Teufel. The mobimed approach to privacy in medical systems. In Proceedings of the 2nd Nordic Workshop on Secure Computer Systems (NORDSEC’97). ????, ????, 1997. To be published.

Capellmann:1996:VBA

[CDFV^{+}96] Carla Capellmann, Ralph Demant, Farhad Fatahi-Vanani, Raphael Galvez-Estrada, Ulrich Nitsche, and Peter Ochsenschläger. Verification by behavior abstraction: A case study of service interaction detection in intelligent telephone networks. In Alur and Henzinger [AH96], pages 466–469. ISBN 3-540-61474-

5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.V47 C38 1996.

Capellmann:1996:CSS

[CDGE^{+}96] Carla Capellmann, Ralph Demant, Raphael Galvez-Estrada, Ulrich Nitsche, and Peter Ochsenschläger. Case study: Service interaction detection by formal verification under behaviour abstraction. In Tiziana Margaria, editor, Proceedings of the 1st International Workshop on Advanced Intelligent Networks (AIN’96), pages 71–90. ????, ????, March 1996.

Demant:1995:ABG

[DFVGE^{+}95] Ralph Demant, Farhad Fatahi-Vanani, Rafael Galvez-Estrada, Ulrich Nitsche, and Peter Ochsenschläger. Abschlußbericht des GMD-/Telekom-Projekts Formale Spezifikationsund Verifikationsmethoden zur Behandlung der Service-Interaction-Problematik — SERVINT. Abschlußbericht, GMD, Bonn, Germany, Dezember 1995.

Demant:1996:ABG

[DFVGE^{+}96a] Ralph Demant, Farhad Fatahi-Vanani, Rafael Galvez-Estrada, Astrid Link, Ulrich Nitsche, Armin Noll, Peter Ochsenschläger, Jürgen Repp, and Roland Rieke. Abschlußbericht des GMD-/Telekom-Projekts Formale Spezifikations- und Verifikationsmethoden zur Behandlung der Service-Interaction-Problematik — SERVINTII. Abschlußbericht, GMD, Bonn, Germany, Dezember 1996.

Demant:1996:ZGT

[DFVGE^{+}96b] Ralph Demant, Farhad Fatahi-Vanani, Rafael Galvez-Estrada, Ulrich Nitsche, and Peter Ochsenschläger. Zwischenbericht des GMD-/Telekom-Projekts Formale Spezifikationsund Verifikationsmethoden zur Behandlung der Service-Interaction-Problematik — SERVINT2. Zwischenbericht, GMD, Bonn, Germany, July 1996.

Dassow:1996:DLT

[DRS96] Jurgen Dassow, Grzegorz Rozenberg, and Arto Salomaa, editors. Developments in language theory II: at the crossroads of mathematics, computer science, and biology : Magdeburg, Germany, 17– 21 July 1995. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2682-9. LCCN QA267.3 .D483 1996.

Morger:1997:DQK

[MNT97a] Othmar Morger, Ulrich Nitsche, and Stephanie Teufel. Datenschutz als Qualitätskriterium für Krankenhausinformationssysteme. In 42. Jahrestagung der Deutschen Gesellschaft für Medizinische Informatik, Biometrie und Epidemologie (GMDS) ’97. ????, ????, 1997.

Morger:1997:MPE

[MNT97b] Othmar Morger, Ulrich Nitsche, and Stephanie Teufel. Mobimed— privacy and efficiency in mobile medical systems (abstract to the poster session). In Proceedings of the 12th International Workshop on Computing in Clinical Laboratories (CCL’97). ????, ????, 1997.

Morger:1997:SCM

[MNT97c] Othmar Morger, Ulrich Nitsche, and Stephanie Teufel. Security concerns for mobile information systems in health care. In Wagner [Wag97], pages 312– 317. ISBN 0-8186-8147-0, 0-8186-8148-9 (case), 0-8186-8149-7 (microfiche). LCCN QA76.9.D3 D15592 1997b.

Morger:1998:PEP

[MNT98] Othmar Morger, Ulrich Nitsche, and Stephanie Teufel. Privacy and efficiency in patient focused health care processes. In Proceedings of the 9th World Congress on Medical Informatics (MedInfo’98). ????, ????, 1998.

Nitsche:1988:EPI

[Nit88]  Ulrich Nitsche.  Erreich-

barkeitsgraphen von Produktnetzen und ihre Auswertung in PROLOG. Arbeitspapiere der GMD 330, Gesellschaft für Mathematik und Datenverarbeitung (GMD), Darmstadt, August 1988.

Nitsche:1991:EEP

[Nit91a] Ulrich Nitsche. Das Erreichbarkeitsproblem für eingeschränkte Produktnetze. Diploma thesis, University of Frankfurt, Germany, February 1991. Presented at the ITHEM-Meeting at the École Normale Supérieur, Paris, 1991.

Nitsche:1991:EP

[Nit91b]  Ulrich Nitsche.  Erreich-

barkeitsanalyse von Produktnetzen. Arbeitspapiere der GMD 521, Gesellschaft für Mathematik und Datenverarbeitung (GMD), Darmstadt, April 1991.

Nitsche:1994:PLT

[Nit94a] U. Nitsche. Propositional linear temporal logic and language homomorphisms. Lecture Notes in Computer Science, 813:265–277, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitsche:1994:SHL

[Nit94b] Ulrich Nitsche. Simple homomorphisms and linear temporal logic. Arbeitspapiere der GMD 889, GMD — Forschungszentrum Informationstechnik, Darmstadt, December 1994.

Nitsche:1994:VMB

[Nit94c] Ulrich Nitsche. A verification method based on homomorphic model abstraction.

In ACM [ACM94], page 393. ISBN 0-89791-654-9. LCCN QA76.9.D5 A33 1994.

Nitsche:1994:VTL

[Nit94d] Ulrich Nitsche. Verifying temporal logic formulas in abstractions of large reachability graphs. In Jörg Desel, Andreas Oberweis, and Wolfgang Reisig, editors, National Workshop on Algorithms and Tools for Petri Nets (Algorithmen und Werkzeuge für Petrinetze, AWP’94), pages 42–47. Humboldt Universität Berlin, Berlin, Germany, 1994.

Nitsche:1996:FLS

[Nit96a] Ulrich Nitsche. A finitary language semantics for propositional linear temporal logic. In Dassow et al. [DRS96], pages 482–489. ISBN 981-02-2682-9. LCCN QA267.3 .D483 1996.

Nitsche:1996:VBA

[Nit96b] Ulrich Nitsche. Verification and behavior abstraction towards a tractable verification technique for large distributed systems. The Journal of Systems and Software, 33(3): 273–285, June 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nitsche:1997:IFC

[Nit97] Ulrich Nitsche. Interpretation of the full computation tree logic CTL^{∗} on sets of infinite sequences. In Adian and Nerode [AN97], pages 276– 282. LCCN QA75.5 .L43 1997.

Nitsche:1998:AFV

[Nit98] Ulrich Nitsche. Application of formal verification and behaviour abstraction to the service interaction problem in intelligent networks. The Journal of Systems and Software, 40(3):227–248, March 1998. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Nitsche:19xx:VCO

[Nitxx] Ulrich Nitsche. Verification of Co-Operating Systems and Behaviour Abstraction. PhD thesis, University of Frankfurt, Germany, 19xx. Handed in in 1996.

Nerode:1994:LFC

[NM94] Anil Nerode and Iu. V. Matiiasevich, editors. Logical foundations of computer science: third international symposium, LFCS ’94, St. Petersburg, Russia, July 11– 14, 1994: proceedings, volume 813 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-58140-5 (Berlin), 0-387-58140-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .L64

Niessner:1997:DRL

[NNO97] Frank Nießner, Ulrich Nitsche, and Peter Ochsenschläger. Deterministic ω-regular liveness properties. In Preproceedings of the 3rd International Conference on Developments in Language Theory (DLT’97). Aristotle University of Thessaloniki, Greece, ????, ????, 1997.

Niessner:1998:DRL

[NNO98] Frank Nießner, Ulrich Nitsche, and Peter Ochsenschläger. Deterministic ω-regular liveness properties. In Proceedings of the 3rd International Conference on Developments in Language Theory (DLT’97). ????, ????, 1998. To be published.

Nitsche:1995:ASP

[NO95a] Ulrich Nitsche and Peter Ochsenschläger. Approximately satisfied properties of systems and simple language homomorphisms. Arbeitspapiere der GMD 965, GMD — Forschungszentrum Informationstechnik, Darmstadt, December 1995.

Nitsche:1995:ZGT

[NO95b] Ulrich Nitsche and Peter Ochsenschläger. Zwischenbericht des GMD-/Telekom-Projekts Formale Spezifikationsund Verifikationsmethoden zur Behandlung der Service-Interaction-Problematik — SERVINT. Zwischenbericht, GMD, Bonn, Germany, Juli 1995.

Nitsche:1996:ASPa

[NO96a] Ulrich Nitsche and Peter Ochsenschläger. Approximately satisfied properties of systems and simple language homomorphisms. In Proceedings of the 1st International Workshop on Formal Methods for Industrial Critical Systems (FMICS’96). ????, ????, March 1996.

Nitsche:1996:ASPb

[NO96b] Ulrich Nitsche and Peter Ochsenschläger. Approximately satisfied properties of systems and simple language homomorphisms. Information Processing Letters, 60(4):201– 206, November 25, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nitsche:1996:SVT

[NOR96] Ulrich Nitsche, Peter Ochsenschläger, and Jürgen Repp. The SH-Verification Tool. Arbeitspapiere der GMD 1007, GMD — Forschungszentrum Informationstechnik, Darmstadt, July 1996.

Nitsche:1997:SVT

[NORR97] Ulrich Nitsche, Peter Ochsenschläger, Jürgen Repp, and Roland Rieke. The SHverification tool. In Proceedings of the 2nd International Workshop on Formal Methods for Industrial Critical Systems (FMICS’97). ????, ????, 1997.

Niessner:2007:CCD

[NUN07] Frank Nießner and Ulrich Ultes-Nitsche. A complete characterization of deterministic regular liveness properties. Theoretical Computer Science, 387(2):187–195, November 12, 2007. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nitsche:1997:RLB

[NW97] Ulrich Nitsche and Pierre Wolper. Relative liveness and behavior abstraction (extended abstract). In ACM [ACM97], pages 45–52. ISBN 0-89791-952-1. ACM order number 536970. Ultes-Nitsche:2002:DWN

[UN02]  Ulrich Ultes-Nitsche.  Do

we need liveness? — approximation of liveness properties by safety properties. Lecture Notes in Computer Science, 2540:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2540/25400279.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2540/25400279.pdf.

Ultes-Nitsche:2007:PSC

[UN07]  Ulrich Ultes-Nitsche.  A

power-set construction for reducing Büchi automata to non-determinism degree two. Information Processing Letters, 101(3):107–111, February 14, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagner:1997:PEI

[Wag97] Roland R. Wagner, editor. Proceedings: Eighth International Workshop on Database and Expert Systems Applications, September 1–2, 1997, Toulouse, France. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8147-0, 0-8186-8148-9 (case), 0-8186-8149-7 (microfiche). LCCN QA76.9.D3 D15592 1997b.